OpenStudy (anonymous):

I need to know the pH (for each) based on the Ka and Kb values given: Acid…………….Ka HCN………4.0*10^-10 HOCl……..2.95*10^-8 Base……………Kb C5H5N…………2.9*10^-9 HONH2…………9.1*10^-9

OpenStudy (anonymous):

@seehearcreate are you good with pH calculations?

OpenStudy (anonymous):

Yep! Okay, so take the negative log of the acid conc. to find the pH. For the bases, since its pOH, you need to do the same but then take that answer and subtract it from 14(:

OpenStudy (anonymous):

do you mind setting one up so I can see? I'm struggling

OpenStudy (anonymous):

for the first one HCN ka = 4.0 * 10^-10 then pH = - log ( ka) for base as @seehearcreate says its pOH then find pOH = -log(Kb) pH + pOH = 14

OpenStudy (anonymous):

ok I got 4.6 is that right @thushananth01

OpenStudy (anonymous):

oh wait hold up

OpenStudy (anonymous):

9.4?? @seehearcreate

OpenStudy (anonymous):

Let me calculate them, hold on. @thushananth01 that's what I was saying about the bases and pOH.

OpenStudy (anonymous):

Is the 5.4 M HOCl pH=7.5

OpenStudy (anonymous):

9.39 7.53 Oh wait, I see what you're saying that it says Kb. Forget about subtracting. Hold on, let me get the bases.

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Ok so I was on the right track. Now for C5H5N is it 5.46

OpenStudy (anonymous):

No, I was wrong, since it's already calculated values for Kb. You'd wouldn't be trying to find the H concentrations but the OH. So, the values are just regular -log(OH-) So it would be 8.53 and 9.04. You'll always know it's not a base if the pH is below 7. That's an acid.

OpenStudy (anonymous):

Ok are those HOCl and HCN? @seehearcreate

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

Ok sweet. HOCl= 8.53 HCN= 9.04 the bases are…? I got C5H5N= pH 5.46 HONH2

OpenStudy (anonymous):

@seehearcreate

OpenStudy (anonymous):

sorry, I just wanna make sure I got this

OpenStudy (anonymous):

Hold on, none of these values are coming back correctly.

OpenStudy (anonymous):

Are you sure you have the acids and bases under the right values? they seem backwards.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Well I'm not sure then. I'm a little busy since I'm trying to finish up my classes before exams tomorrow, so I'll try to get someone else for you. @abb0t

OpenStudy (anonymous):

no worries I completely understand, thank you! @seehearcreate